OpenStudy (camzzzie):

PLEASE HELP I know how you are supposed to do it but I cant understand this one:( Solve the equation. Check for extraneous solutions. 9|9 – 8x| = 2x + 3

OpenStudy (camzzzie):

Divide by 9 9-8x= 2x+ 1/3 or 9-8x= -(2x+ 1/3) Add 8x both sides Minus 1/3 both sides

OpenStudy (camzzzie):

thats how im trying to do it but idk if thats what the nine needs to do

OpenStudy (phi):

when you divide by 9 on both sides , the right side becomes \[ \frac{2x +3}{9} \] which can also be written as \[ \frac{2}{9}x + \frac{1}{3} \]

OpenStudy (phi):

the rest looks ok. For the positive equation, we would have \[ 9 - 8x = \frac{2}{9}x + \frac{1}{3} \\ 9 = 8x+\frac{2}{9}x + \frac{1}{3} \\ 9- \frac{1}{3} = 8x+\frac{2}{9}x \] after combining the fractions \[ \frac{26}{3} = \frac{74}{9}x \] and \[ x = \frac{\cancel{9}3}{\cancel{74}37} \cdot \frac{\cancel{26}13}{\cancel{3}} =\frac{39}{37}\]

OpenStudy (camzzzie):

@phi thanks:)